OpenStudy (123ab456c):

Which equation can be used to find the answer? After two months, Rachel has saved enough money to buy an MP3 player. During the second month, she saved $87 and she pays $199 for the MP3 player. How much did Rachel save during the first month? A. m = 87 • 2 m = 174 $174 B. 199 + 87 = m m = 286 $286 C. m + 87 = 199 m = 112 $112 D. m – 87 = 199 m = 287 $287

OpenStudy (123ab456c):

if you help me i will medal

OpenStudy (shiburinga):

What do you think it is?

OpenStudy (123ab456c):

thats why im asking you to tell me the answer:)

OpenStudy (123ab456c):

are you smart becuase youre smartness says 57

OpenStudy (shiburinga):

Sorry, not gonna tell you but help you find it. :P Well, the amount saved in first month (m) is uknown. The second month, shae save about $87 dollars. the equation is m+87 for now.

OpenStudy (123ab456c):

then DONT help me

OpenStudy (shiburinga):

Um.. I'm helping, just not giving you the answer.

OpenStudy (123ab456c):

i want pooja195 to helpme becuse her smart is 99

OpenStudy (shiburinga):

I already got warned for the giving someone the answer and not showing them now how to do it. And, I now now just giving the answer isn't going to help anyone. And Pooja is going to show you how to get the answer, too.

OpenStudy (123ab456c):

then how come everyone else tells me the answer!!! :C

OpenStudy (shiburinga):

I have no idea.. and what my point is, from my understanding, we are supposed to teach other people what they don't.

OpenStudy (shiburinga):

Maybe they were like I was... finding the easiest way out of the question, trying to get medals and stuff easier. :P

pooja195 (pooja195):

\(\rm\color{green}{~~~~~~~~~~~~~~~Learning}\color{chocolate}{~is~alike~to~the~journey~of~an~}\color{green}{adventure.}\) \(\rm\color{chocolate}{~~~~An~adventure~with~the~}\color{green}{goal}\color{chocolate}{~of~attaining~a~greater~}\color{gold}{knowledge}\color{chocolate}{.}\) \(\rm\color{chocolate}{However,~it~is~}\color{darkred}{not~only}\color{chocolate}{~the~knowledge~alone~that~is~of~importance,}\) \(\rm\color{chocolate}{~~~~~For~it~is~often~times~our~}\color{green}{journey}\color{chocolate}{~that}\color{lime}{~inspires~}\color{chocolate}{our~}\color{gold}{knowledge.~whitemonsterbunny17}\)

OpenStudy (shiburinga):

#pooja195 #LaTex #Whitemonsternbunny17 :P

OpenStudy (123ab456c):

please just tell me the answer i understand but just need the answer! pleassssseeeeee

OpenStudy (shiburinga):

Just read through this answer.. :P http://ec2-52-0-184-118.compute-1.amazonaws.com/updates/53925126e4b0c4e31004441e And you sure you understand?

OpenStudy (123ab456c):

Which equation can be used to find the answer? After two months, Rachel has saved enough money to buy an MP3 player. During the second month, she saved $87 and she pays $199 for the MP3 player. How much did Rachel save during the first month? A. m = 87 • 2 m = 174 $174 B. 199 + 87 = m m = 286 $286 C. m + 87 = 199 m = 112 $112 D. m – 87 = 199 m = 287 $287

OpenStudy (123ab456c):

JUST TELL ME THE ANSWER YOU IDIOT!!!!!

OpenStudy (shiburinga):

The answer is on that site I gave you.

OpenStudy (123ab456c):

SO the answer is C am i correct

OpenStudy (shiburinga):

Yes.

OpenStudy (123ab456c):

thank you and sorry for being mean

OpenStudy (123ab456c):

you where just helping

OpenStudy (shiburinga):

It's cool, everyone gets aggrevated at one point. :)

OpenStudy (123ab456c):

ok:):):):):):)

OpenStudy (shiburinga):

Good luck on this lesson. ^.^